Find the missing measurements of each rectangle if area = 216 and perimeter = 66
Answer by stanbon(75887)

Find the missing measurements of each rectangle if
area = 216 and perimeter = 66
---------------
L + W = 33
L*W = 216
----
Substitute for "L" and solve for "W"::
(33-W)*W = 216
-----
W^2 - 33W + 216 = 0
(W-9)(W-24) = 0
---
Width = 9, then Length = 24
OR
Width = 24, then Length = 9
-----------------
Cheers,
Stan H.
-----------------